OpenStudy (anonymous):

Easy question: How do you make this equation equal to y, as in the equation of a line? x=sqrroot (4-y^2)

OpenStudy (anonymous):

how did the y loose its square?

OpenStudy (anonymous):

\[x=\sqrt{4-y^2}\] \[x^2=4-y^2\] \[x^2+y^2=4\]

OpenStudy (anonymous):

it is the right half of a circle with center and the origin and radius 2

OpenStudy (anonymous):

if you want to solve for y be careful start with \[x^2+y^2=4\] \[y^2=4-x^2\] \[y=\pm\sqrt{4-x^2}\]

OpenStudy (anonymous):

thanks! just one question though. By right half do you mean quadrants 1 and 4?

OpenStudy (anonymous):

the equation for the circle with center (0,0) and radius 2 is \[x^2+y^2=4\] if you solve this for x you actually get \[x=\pm\sqrt{4-y^2}\]

OpenStudy (anonymous):

note the important "plus or minus" because of course you do not know which

OpenStudy (anonymous):

if you just write \[x=\sqrt{4-y^2}\]without the plus or minus, that means you are forcing x to be positive, because that is what the radical sign means. so if x is positive it is the right half of the circle. quadrants one and 4 yes.

OpenStudy (anonymous):

if you write \[y=\sqrt{4-x^2}\] you get the upper half of the circle

OpenStudy (anonymous):

so should I graph the upper or right side?